The problem is to maximise utility u(x,y) = 2*x +y s.t. x,y≥0 and p*x + q*y ≤w, where p=17.6 and q=3.4 and w=1. The solution to this problem is denoted (x_0, y_0) = (x(p, q, w), y(p, q, w)). The solution is the global max.

Find du(x_0,y_0)/dp evaluated at the parameters (p, q, w) = (17.6, 3.4, 1). Write the answer as a number in decimal notation with at least two digits after the decimal point.
Answer #1

st solution objective is to maximise u(x, y) = 2x+y i ) px+qy sw 4) where p=17-6 and d Y = 3.4 this problem graphically by cofrom figot a A q - 0294 icz 9 T ico ic, Ві -0.0568 line Budget figil. The Global optimum the cetility y a 18 دا where from tů so r. O Too 0.294 q becaus at optimum u(x0, y.o) independent of price of a p. _ = 0 , 1 u (20.7.0) 0.0.294 du(xo, Y.o dh an